Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis

Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 4,400 units of product were as follows:

Standard Costs Actual Costs
Direct materials 5,700 lb. at $5.10 5,600 lb. at $5.00
Direct labor 1,100 hrs. at $18.00 1,130 hrs. at $18.20
Factory overhead Rates per direct labor hr.,
based on 100% of normal
capacity of 1,150 direct
labor hrs.:
Variable cost, $3.50 $3,810 variable cost
Fixed cost, $5.50 $6,325 fixed cost

Each unit requires 0.25 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct materials price variance $
Direct materials quantity variance
Total direct materials cost variance $

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct labor rate variance $
Direct labor time variance
Total direct labor cost variance $

c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance $
Fixed factory overhead volume variance
Total factory overhead cost variance
Answer #1

a. Material price variance = (5.10-5)*5600 = 560 favorable

Material quantity variance = (5700-5600)*5.10 = 510 favorable

Total material variance = 5700*5.10-5600*5 = $1070 favorable

b. Labor rate variance = (18-18.20)*1130 = 226 unfavorable

Labor time variance = (1100-1130)*18 = 540 unfavorable

Total labor cost variance = 1100*18-1130.18.20 = 766 unfavorable